OpenStudy (anonymous):

I'm looking for an equation that will help me discover the original cost of an appliance before the discount was applied. Discount is 20% and the cost after is 467.20. I would rather not get the answer to the problem but just help setting up the equation. I did one here lastnight with lots of help but I am still struggling. ty in advance for any help, Denise

OpenStudy (anonymous):

x - 0.20x = 467.20 where x is the original cost

OpenStudy (anonymous):

new price = full price - (full price * discount amount) 467.20 = x - (x * .20) 467.20 = x -.20x 467.20 = .80x 467.20/.80 = x 584 = x

OpenStudy (anonymous):

Well thank you both so much. I hope I can help someone else soon. This forum is something I would have liked to have found 2 terms ago, LOL!! Denise

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

Harkirat, why is there an x with the 0.20? That's the only part I don't understand? Are you still here?

OpenStudy (anonymous):

yes...i am here....

OpenStudy (anonymous):

u have to calculate 20% of x and subtract it from x to get final price/cost... Is it clear now??

OpenStudy (anonymous):

Thank you Juanita, very much, Denise:)

OpenStudy (anonymous):

Yes, I believe so, I am going to work on it now:) ty again:) Denise

OpenStudy (anonymous):

Juanita, may I bother you one more time? How did you get the .80x?

OpenStudy (anonymous):

x - 0.20x = (1 - 0.20) x = 0.80x

OpenStudy (anonymous):

Ok, thank you harkirat, I will try again to calculate it on my own. Denise:)

OpenStudy (anonymous):

Ok, so you are taking 1 as a whole number and subtracting .20 to get .80. I see that. I will figure it out;) I hope, LOL!!

OpenStudy (anonymous):

x means 1x

OpenStudy (anonymous):

You are figuring the X as 1x right?

OpenStudy (anonymous):

suppose u hv $ 100 20% of it = (20/100) * 100 = $20 left is $ 80 i.e. 80 out of 100 i.e. 80/100 = 0.80

OpenStudy (anonymous):

Ok, I gotcha now:) ty again, especially for your patience:) denise

OpenStudy (anonymous):

welcome..☺

OpenStudy (anonymous):

I got it, I mean I figured out how you guys came to that and I learned how to make the equation and work it:) Ty sooooooooo much, Denise